Graphing Using The Slope-Intercept Form
please show the solutions of the ff. equations:
1)5x=2y-10
2)4x-3y=12
3)2x+3y=12
Pls. reply as soon as as possible.

Answer by mananth(16946)

Graphing Using The Slope-Intercept Form
please show the solutions of the ff. equations:
1)5x=2y-10
2y= 5x+10
y = (5x/2)+5
m= slope = 5/2
find the intercepting points on the x & y axis
..
x=0 y=5 (0,5)
y=0, x=-2 (-2,0)
plot these points join the points to get the line graph.
...

I have walked you through with the first. the others are same.
2)4x-3y=12
3)2x+3y=12
m.ananth@hotmail.ca